# Unveiling the Role of Learning Rate Schedules via Functional Scaling Laws

### v1

@@ C.4 The Case of Top-M Features
$\mathbb{E}[\mathcal{E}_{t}]$ $\eqsim g_{s}(t)+M^{-s\beta}+\sigma^{2}\int_{0}^{t}\mathcal{K}(t-\tau)\gamma_{\tau}\mathop{}\!\mathrm{d}\tau$
$\qquad\quad+\int_{0}^{t}\mathcal{K}(t-\tau)\gamma_{\tau}\left(g_{s}(\tau)+M^{-s\beta}+\sigma^{2}\int_{0}^{\tau}\mathcal{K}(\tau-r)\gamma_{r}\mathop{}\!\mathrm{d}r\right)\mathop{}\!\mathrm{d}\tau$
$=g_{s}(t)+M^{-s\beta}+(\sigma^{2}+M^{-s\beta})p(t)+q(t)+\sigma^{2}\int_{0}^{t}\mathcal{K}(t-\tau)\gamma_{\tau}\int_{0}^{\tau}\mathcal{K}(\tau-r)\gamma_{r}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}\tau.$
The last term can be bounded as follows
$\sigma^{2}\int_{0}^{t}\mathcal{K}(t-\tau)$ $\gamma_{\tau}\int_{0}^{\tau}\mathcal{K}(\tau-r)\gamma_{r}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}\tau$
$=\sigma^{2}\int_{0}^{t}\gamma_{r}\int_{r}^{t}\mathcal{K}(t-\tau)\mathcal{K}(\tau-r)\gamma_{\tau}\mathop{}\!\mathrm{d}\tau\mathop{}\!\mathrm{d}r$
$\leqslant\sigma^{2}\int_{0}^{t}\gamma_{r}\int_{r}^{t}\mathcal{K}(t-\tau)\mathcal{K}(\tau-r)\gamma_{\max}\mathop{}\!\mathrm{d}\tau\mathop{}\!\mathrm{d}r$
$\stackrel{{\scriptstyle(i)}}{{\leqslant}}C\gamma_{\max}\sigma^{2}\int_{0}^{t}\gamma_{r}\mathcal{K}(t-r)\mathop{}\!\mathrm{d}r=C\gamma_{\max}\sigma^{2}p(t),$
where $(i)$ uses Lemma C.9. Suppose $C\gamma_{\max}<\frac{1}{2}$, we have
$\mathbb{E}[\mathcal{E}_{t}]\eqsim g_{s}(t)+M^{-s\beta}+(\sigma^{2}+M^{-s\beta})p(t)+q(t).$
Suppose $\gamma_{\max}$ to be sufficiently small such that
$p(t)\leqslant\gamma_{\max}\int_{0}^{t}g_{2-1/\beta}(\tau)\mathop{}\!\mathrm{d}\tau\leqslant\gamma_{\max}\frac{\beta}{2(\beta-1)}<\frac{1}{2}.$
Then, we have
$\mathbb{E}[\mathcal{E}_{t}]$ $\eqsim g_{s}(t)+M^{-s\beta}+\sigma^{2}p(t)+q(t)$
$=g_{s}(t)+M^{-s\beta}+\int_{0}^{t}\mathcal{K}(t-\tau)(\sigma^{2}+g_{s}(\tau))\gamma_{\tau}\mathop{}\!\mathrm{d}\tau.$
The case of $M^{\beta}\leqslant t$:
The condition of Lemma C.9 does not hold. In this case, we first analyze the left-hand side of the equation. The key idea is to make a small modification to $B(t)$ and use similar methods to prove the result.
By (C.4),
$B(t)=\int_{0}^{t}\mathcal{K}(t-\tau)\,c_{\tau}\mathbb{E}[\mathcal{E}_{\tau}]\gamma_{\tau}\mathop{}\!\mathrm{d}\tau\leqslant\int_{0}^{t}G_{2-1/\beta}(t-\tau)\,c_{\tau}\mathbb{E}[\mathcal{E}_{\tau}]\gamma_{\tau}\mathop{}\!\mathrm{d}\tau=:\tilde{B}(t).$
We next proceed to estimate $\tilde{B}(t)$ in an analogous manner.
$\tilde{B}(t)$ $=\int_{0}^{t}G_{2-1/\beta}(t-\tau)\,c_{\tau}\mathbb{E}[\mathcal{E}_{\tau}]\gamma_{\tau}\mathop{}\!\mathrm{d}\tau$
$=\int_{0}^{t}G_{2-1/\beta}(t-\tau)\,c_{\tau}\gamma_{\tau}\left(g_{s}(\tau)+M^{-s\beta}+\sigma^{2}\int_{0}^{\tau}\mathcal{K}(\tau-r)c_{r}\gamma_{r}\mathop{}\!\mathrm{d}r\right)\mathop{}\!\mathrm{d}\tau$
$\qquad\quad+\int_{0}^{t}G_{2-1/\beta}(t-\tau)\,c_{\tau}\gamma_{\tau}\int_{0}^{\tau}\mathcal{K}(\tau-r)c_{r}\mathbb{E}[\mathcal{E}_{r}]\gamma_{r}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}\tau.$
Then we calculate the last term. By Fubini’s theorem,
$\int_{0}^{t}G_{2-1/\beta}(t-\tau)\,c_{\tau}\gamma_{\tau}$ $\int_{0}^{\tau}\mathcal{K}(\tau-r)c_{r}\mathbb{E}[\mathcal{E}_{r}]\gamma_{r}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}\tau$
$=\int_{0}^{t}c_{r}\mathbb{E}[\mathcal{E}_{r}]\gamma_{r}\int_{r}^{t}G_{2-1/\beta}(t-\tau)\mathcal{K}(\tau-r)\gamma_{\tau}\mathop{}\!\mathrm{d}\tau\mathop{}\!\mathrm{d}r$
$\leqslant\int_{0}^{t}c_{r}\mathbb{E}[\mathcal{E}_{r}]\gamma_{r}\int_{r}^{t}G_{2-1/\beta}(t-\tau)G_{2-1/\beta}(\tau-r)\gamma_{\max}\mathop{}\!\mathrm{d}\tau\mathop{}\!\mathrm{d}r$
$\leqslant C\gamma_{\max}\int_{0}^{t}c_{r}\mathbb{E}[\mathcal{E}_{r}]\gamma_{r}G_{2-1/\beta}(t-r)\mathop{}\!\mathrm{d}r$
$=C\gamma_{\max}\tilde{B}(t),$
where the last inequality is by Lemma C.10. Therefore, we get
$\tilde{B}(t)\leqslant\frac{1}{1-C\gamma_{\max}}\int_{0}^{t}G_{2-1/\beta}(t-\tau)\,c_{\tau}\gamma_{\tau}\left(g_{s}(\tau)+M^{-s\beta}+\sigma^{2}\int_{0}^{\tau}\mathcal{K}(\tau-r)c_{r}\gamma_{r}\mathop{}\!\mathrm{d}r\right)\mathop{}\!\mathrm{d}\tau.$
$\tilde{B}(t)\geqslant\int_{0}^{t}G_{2-1/\beta}(t-\tau)\,c_{\tau}\gamma_{\tau}\left(g_{s}(\tau)+M^{-s\beta}+\sigma^{2}\int_{0}^{\tau}\mathcal{K}(\tau-r)c_{r}\gamma_{r}\mathop{}\!\mathrm{d}r\right)\mathop{}\!\mathrm{d}\tau.$
As $C\gamma_{\max}\leqslant\frac{1}{2}$, we have
$\tilde{B}(t)\eqsim\int_{0}^{t}G_{2-1/\beta}(t-\tau)\,c_{\tau}\gamma_{\tau}\left(g_{s}(\tau)+M^{-s\beta}+\sigma^{2}\int_{0}^{\tau}\mathcal{K}(\tau-r)c_{r}\gamma_{r}\mathop{}\!\mathrm{d}r\right)\mathop{}\!\mathrm{d}\tau.$
Next we bound the last term
$\int_{0}^{t}$ $G_{2-1/\beta}(t-\tau)\gamma_{\tau}\int_{0}^{\tau}\mathcal{K}(\tau-r)\gamma_{r}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}\tau$
$=\int_{0}^{t}\gamma_{r}\int_{r}^{t}G_{2-1/\beta}(t-\tau)\mathcal{K}(\tau-r)\gamma_{\tau}\mathop{}\!\mathrm{d}\tau\mathop{}\!\mathrm{d}r$
$\leqslant\int_{0}^{t}\gamma_{r}\int_{r}^{t}G_{2-1/\beta}(t-\tau)G_{2-1/\beta}(\tau-r)\gamma_{\max}\mathop{}\!\mathrm{d}\tau\mathop{}\!\mathrm{d}r$
$\eqsim\gamma_{\max}\int_{0}^{t}G_{2-1/\beta}(t-r)\gamma_{r}\mathop{}\!\mathrm{d}r.$
where the last inequality is again by Lemma C.10.
Also we have
$\int_{0}^{t}\mathcal{K}(\tau)\mathop{}\!\mathrm{d}\tau$ $=\int_{0}^{t}\int_{M^{-\beta}}^{1}u^{1-1/\beta}e^{-2u\tau}\mathop{}\!\mathrm{d}u\mathop{}\!\mathrm{d}\tau$
$=\int_{M^{-\beta}}^{1}u^{1-1/\beta}\frac{1-e^{-2ut}}{2u}\mathop{}\!\mathrm{d}u$
$\leqslant\int_{M^{-\beta}}^{1}u^{1-1/\beta}\frac{1}{2u}\eqsim 1,$
Since we assumed that $t$ is sufficiently large, we also have $\int_{0}^{t}\mathcal{K}(\tau)\mathop{}\!\mathrm{d}\tau\eqsim 1$, and similarly
$\int_{0}^{t}(G_{2-1/\beta}(\tau)-\mathcal{K}(\tau))\gamma_{\tau}\mathop{}\!\mathrm{d}\tau\lesssim\gamma_{\max}\int_{0}^{M^{-\beta}}u^{-1/\beta}\mathop{}\!\mathrm{d}u\eqsim\gamma_{\max}M^{1-\beta},$
which implies $\int_{0}^{t}G_{2-1/\beta}(\tau)\gamma_{\tau}\mathop{}\!\mathrm{d}\tau\eqsim\int_{0}^{t}\mathcal{K}(\tau)\gamma_{\tau}\mathop{}\!\mathrm{d}\tau$.
Therefore, continuing the computation of $\tilde{B}(t)$, since
$M^{\beta}<t\implies M^{-2\beta+1}\geqslant t^{-(2-\frac{1}{\beta})}\gtrsim G_{2-1/\beta}(t)\geqslant\mathcal{K}(t),$
we have $M^{-\beta}+G_{2-1/\beta}(t)\eqsim M^{-\beta}+\mathcal{K}(t)$, hence similar to the proof in the first case, let
$p(t):=\int_{0}^{t}G_{2-1/\beta}(\tau)\gamma_{\tau}\mathop{}\!\mathrm{d}\tau\lesssim\gamma_{\max},\quad q(t):=\int_{0}^{t}G_{2-1/\beta}(t-\tau)g_{s}(\tau)\gamma_{\tau}\mathop{}\!\mathrm{d}\tau.$
Since $\int_{0}^{t}g_{s}(\tau)\mathop{}\!\mathrm{d}\tau\lesssim\max\{1,M^{-s\beta+\beta}\}$, by $s\leqslant 2-1/\beta$, we have $\min(M^{-s\beta},M^{-\beta})\geqslant M^{-2\beta+1}$, we have
$q(t)+M^{-s\beta}$ $\eqsim\int_{0}^{t}(G_{2-1/\beta}(t-\tau)+M^{-2\beta+1})g_{s}(\tau)\gamma_{\tau}\mathop{}\!\mathrm{d}\tau+M^{-s\beta}$
$\eqsim\int_{0}^{t}(g_{2-1/\beta}(t-\tau)+M^{-2\beta+1})g_{s}(\tau)\gamma_{\tau}\mathop{}\!\mathrm{d}\tau+M^{-s\beta}$
$\eqsim\int_{0}^{t}\mathcal{K}(t-\tau)g_{s}(\tau)\gamma_{\tau}\mathop{}\!\mathrm{d}\tau+M^{-s\beta}.$
Therefore by $p(t)\eqsim\int_{0}^{t}\mathcal{K}(\tau)\mathop{}\!\mathrm{d}\tau\lesssim\gamma_{\max}$, we get
$\tilde{B}(t)$ $\eqsim\int_{0}^{t}G_{2-1/\beta}(t-\tau)\,c_{\tau}\gamma_{\tau}\left(g_{s}(\tau)+M^{-s\beta}+\sigma^{2}\int_{0}^{\tau}\mathcal{K}(\tau-r)c_{r}\gamma_{r}\mathop{}\!\mathrm{d}r\right)\mathop{}\!\mathrm{d}\tau$
$\eqsim q(t)+M^{-s\beta}p(t)+\gamma_{\max}\sigma^{2}p(t)$
$\lesssim q(t)+\gamma_{\max}M^{-s\beta}+\gamma_{\max}\sigma^{2}\int_{0}^{t}\mathcal{K}(t-\tau)\gamma_{\tau}\mathop{}\!\mathrm{d}\tau.$
Substituting this into the expression of $\mathcal{E}_{t}$ (Lemma C.12), note that $\mathcal{K}(t)=g_{2-1/\beta}(t)$, we have
$\mathbb{E}[\mathcal{E}_{t}]$ $\lesssim g_{s}(t)+\tilde{B}(t)+M^{-s\beta}+\sigma^{2}\int_{0}^{t}\mathcal{K}(t-\tau)\gamma_{\tau}\mathop{}\!\mathrm{d}\tau$
$\lesssim g_{s}(t)+q(t)+M^{-s\beta}+\sigma^{2}\int_{0}^{t}\mathcal{K}(t-\tau)\gamma_{\tau}\mathop{}\!\mathrm{d}\tau$
$\lesssim M^{-s\beta}+g_{s}(t)+\int_{0}^{t}\mathcal{K}(t-\tau)(\sigma^{2}+g_{s}(t))\gamma_{\tau}\mathop{}\!\mathrm{d}\tau.$
On the other hand, clearly $\mathbb{E}[\mathcal{E}_{t}]\gtrsim g_{s}(t)$, thus by Lemma C.12,
$\mathbb{E}[\mathcal{E}_{t}]$ $\eqsim M^{-s\beta}+g_{s}(t)+\int_{0}^{t}\mathcal{K}(t-\tau)(\sigma^{2}+\mathbb{E}[\mathcal{E}_{\tau}])\gamma_{\tau}\mathop{}\!\mathrm{d}\tau$
$\gtrsim M^{-s\beta}+g_{s}(t)+\int_{0}^{t}\mathcal{K}(t-\tau)(\sigma^{2}+g_{s}(t))\gamma_{\tau}\mathop{}\!\mathrm{d}\tau,$
which completes the proof. ∎
Corollary C.14.
Suppose that $\gamma_{\max}:=\sup_{t}\gamma_{t}$ is sufficiently small and $M=\infty$. Then, for $s\in(0,+\infty)$ we have
$\mathbb{E}[\mathcal{E}_{t}]\eqsim G_{s}(t)+\int_{0}^{t}\mathcal{K}(t-\tau)\left(\sigma^{2}+G_{s}(\tau)\right)\gamma_{\tau}\mathop{}\!\mathrm{d}\tau.$
Proof.
By the case $M^{\beta}>t$ in the proof of Proposition C.13, the result follows. Note that in the case $M^{\beta}>t$, we don’t need the condition of $s\leqslant 2-1/\beta$. ∎
C.5 The Case of Random Features
In this section, we prove Theorem 4.1 for the case of random features. The proof proceed similarly as the top-$M$ case. First we introduce some concentration properties of the random martix $\mathbf{W}\mathbf{H}\mathbf{W}^{\top}$ from [39], then we estimate the upper and lower bound of each term of the Volterra equation in Proposition C.5.
C.5.1 Concentration Inequalities
@@
There exist $\beta$-dependent constants $0<c_{1}<c_{2}$ such that it holds with probability at least $1-e^{-\Omega(M)}$ for all $j\in[M]$ that
$c_{1}j^{-\beta}\leqslant\mu_{j}(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})\leqslant c_{2}j^{-\beta}$
@@
C.5.2 Upper and Lower Bounds
Let $\hat{\lambda}_{j}=\mu_{j}(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})$.
@@ C.5.2 Upper and Lower Bounds
With probability at least $1-e^{-\Omega(M)}$, we have
$\mathrm{tr}(\mathbf{S}\mathbf{A}^{\top}_{t-\tau}\mathbf{H}\mathbf{A}_{t-\tau}\mathbf{S})=\sum_{j=1}^{M}e^{-2(t-\tau)\hat{\lambda}_{j}}\hat{\lambda}_{j}^{2}\eqsim\mathcal{K}(t-\tau).$
@@
$=\sum_{a,b=0}^{\infty}\frac{1}{a!b!}(-t+\tau)^{a+b}\sum_{j=1}^{M}\hat{\lambda}_{j}^{a+b+2}$
$=\sum_{j=1}^{M}e^{-2(t-\tau)\hat{\lambda}_{j}}\hat{\lambda}_{j}^{2}.$
By Lemma C.15, $\hat{\lambda}_{j}\eqsim\lambda_{j}\eqsim j^{-\beta}$, hence the summation can be estimated by the integral
$\sum_{j=1}^{M}e^{-2(t-\tau)\hat{\lambda}_{j}}\hat{\lambda}_{j}^{2}\eqsim\int_{1}^{M}e^{-2(t-\tau)x^{-\beta}}x^{-2\beta}\mathop{}\!\mathrm{d}x=\int_{M^{-\beta}}^{1}e^{-2(t-\tau)u}u^{2-\frac{1}{\beta}}\mathop{}\!\mathrm{d}u=\mathcal{K}(t-\tau),$
where the last but second equality is a change of variable $u=x^{-\beta}$ in the integral. ∎
@@
With probability at least $1-e^{-\Omega(M)}$, we have
$\mathbf{u}_{0}^{\top}\mathbf{A}^{\top}_{t}\mathbf{H}\mathbf{A}_{t}\mathbf{u}_{0}\lesssim M^{-s\beta}+g_{s}(t).$
@@
$P_{t}(x)=\sum_{a+b\geqslant 0}\frac{1}{a!b!}(-t)^{a+b}x^{a+b+1}.$
Note that when $x\in\mathbb{R}$, we have $P_{t}(x)=xe^{-2tx}$, hence the eigenvalues of $\mathbf{M}_{t}$ is exactly $P_{t}(\hat{\lambda}_{j})$. Since
@@
Then by Lemma C.21, we can derive an upper bound. Since $s<1$,
@@
$\eqsim\frac{1}{t}\sum_{j=1}^{k}j^{-1-s(\beta-1)}+\sum_{j=k+1}^{N}j^{-1-s\beta}\eqsim\frac{k^{-s(\beta-1)}}{t}+k^{-s\beta}.$
@@
$\mathbf{u}_{0}^{\top}\mathbf{A}^{\top}_{t}\mathbf{H}\mathbf{A}_{t}\mathbf{u}_{0}\lesssim\max\{t^{-s},M^{-s\beta}\}.$
Now when $M^{\beta}<t$, we have $g_{s}(t)\eqsim\frac{1}{t^{s}}$ by Lemma C.7, the conclusion follows. ∎
@@
It holds with probability at least $1-e^{-\Omega(M)}$ that
$\mathbf{u}_{0}^{\top}\mathbf{A}^{\top}_{t}\mathbf{H}\mathbf{A}_{t}\mathbf{u}_{0}\gtrsim\max\{g_{s}(t),M^{-s\beta}\}.$
@@
where the last inequality is by Von Neumann’s trace inequality. Note that $\mathbf{M}_{t}=P_{t}(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})$, we then get
@@
$=\sum_{i=1}^{M}e^{-2t\hat{\lambda}_{i}}\hat{\lambda}_{i}^{-1}\mu_{i}\left(\mathbf{W}\mathbf{H}\mathbf{u}_{0}\mathbf{u}_{0}^{\top}\mathbf{H}\mathbf{W}^{\top}\right).$
Note that $\mathbf{u}_{0}=\mathbf{v}^{*}$, by Assumption 2.2 and 2.3,
$\mathbf{u}_{0}^{\top}\mathbf{A}^{\top}_{t}\mathbf{H}\mathbf{A}_{t}\mathbf{u}_{0}$ $\geqslant\sum_{i=1}^{M}e^{-2t\hat{\lambda}_{i}}\hat{\lambda}_{i}^{-1}\mu_{i}\left(\mathbf{W}\mathbf{H}\mathbf{u}_{0}\mathbf{u}_{0}^{\top}\mathbf{H}\mathbf{W}^{\top}\right)$
$\eqsim\sum_{i=1}^{M}e^{-2t\hat{\lambda}_{i}}\hat{\lambda}_{i}^{-1}\mu_{i}(\mathbf{W}\mathbf{H}^{1+s+1/\beta}\mathbf{W}^{\top})$
$\gtrsim\sum_{i=1}^{M}e^{-2t\hat{\lambda}_{i}}\hat{\lambda}_{i}^{s+1/\beta}$
$\eqsim\int_{1}^{M}e^{-2tx^{-\beta}}x^{-1-s\beta}\mathop{}\!\mathrm{d}x$
$=\int_{M^{-\beta}}^{1}e^{-2tu}u^{s-1}\mathop{}\!\mathrm{d}u=g_{s}(t).$
Here we used Lemma C.15.
@@
$\mathbf{u}_{0}^{\top}\mathbf{A}^{\top}_{t}\mathbf{H}\mathbf{A}_{t}\mathbf{u}_{0}\geqslant\|(\textbf{I}-\mathbf{H}^{\frac{1}{2}}\mathbf{W}^{\top}(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})^{-1}\mathbf{W}\mathbf{H}^{\frac{1}{2}})\mathbf{H}^{\frac{1}{2}}\mathbf{u}_{0}\|^{2}$

### v3

@@ C.4 The Case of Random-$M$ Features
$\mathbb{E}[\mathcal{E}_{t}]\;\eqsim\;M^{-s\beta}+\widehat{e}_{M}(t)+\int_{0}^{t}\widehat{\mathcal{K}}_{M}(t-\tau)\,\gamma_{\tau}\bigl(\mathbb{E}[\mathcal{E}_{\tau}]+\sigma^{2}\bigr)\,\mathop{}\!\mathrm{d}\tau,$ (42)
where
$\widehat{e}_{M}(t):=\sum_{j=1}^{M}\widehat{\lambda}_{j}|\theta^{*}_{j}|^{2}e^{-2\widehat{\lambda}_{j}t},\quad\widehat{\mathcal{K}}_{M}:=\sum_{j=1}^{M}\widehat{\lambda}_{j}^{2}e^{-2\widehat{\lambda}_{j}t},$
and $\widehat{\lambda}_{j}:=\mu_{j}(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})$ is the $j$-th eigenvalue of the random matrix.
Theorem 4.6 then follows by applying exactly the same argument as in the top-$M$ case. It therefore remains to prove Proposition C.11. The key idea is to show that the spectrum of the random matrix $\mathbf{W}\mathbf{H}\mathbf{W}^{\top}\in\mathbb{R}^{M\times M}$ closely matches that of the top-$M$ truncation, namely,
$\widehat{\lambda}_{j}:=\mu_{j}(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})\;\eqsim\;\lambda_{j},\qquad\forall\;1\leqslant j\leqslant M.$
C.4.1 Concentration Inequalities
@@
For $\mathbf{H}=\operatorname{diag}\{\lambda_{1},\lambda_{2},\dots,\lambda_{N}\}$ such that $\lambda_{j}\eqsim j^{-\beta}$ where $\beta>1$, there exists $\beta$-dependent constants $0<c_{1}<c_{2}$ such that it holds with probability at least $1-e^{-\Omega(M)}$ for all $j\in[M]$ that
$c_{1}\lambda_{j}\leqslant\mu_{j}(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})\leqslant c_{2}\lambda_{j}$
@@
C.4.2 Upper and Lower Bounds
Let $\widehat{\lambda}_{j}=\mu_{j}(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})$.
@@ C.4.2 Upper and Lower Bounds
With probability at least $1-e^{-\Omega(M)}$, for $s>0$ we have
$\mathrm{tr}(\mathbf{S}\mathbf{A}^{\top}_{t-\tau}\mathbf{H}\mathbf{A}_{t-\tau}\mathbf{S})=\sum_{j=1}^{M}e^{-2(t-\tau)\widehat{\lambda}_{j}}\widehat{\lambda}_{j}^{2}=:\widehat{\mathcal{K}}_{M}(t-\tau).$
@@
$=\sum_{a,b=0}^{\infty}\frac{1}{a!b!}(-t+\tau)^{a+b}\sum_{j=1}^{M}\widehat{\lambda}_{j}^{a+b+2}$
$=\sum_{j=1}^{M}e^{-2(t-\tau)\widehat{\lambda}_{j}}\widehat{\lambda}_{j}^{2},$
which completes the proof. ∎
@@
With probability at least $1-e^{-\Omega(M)}$, for $0<s\leqslant 1$ we have
$\mathbf{u}_{0}^{\top}\mathbf{A}^{\top}_{t}\mathbf{H}\mathbf{A}_{t}\mathbf{u}_{0}\lesssim M^{-s\beta}+\widehat{e}_{M}(t).$
@@
$P_{t}(x)$ $:=\sum_{a+b\geqslant 0}\frac{1}{a!b!}(-t)^{a+b}x^{a+b+1}.$
Note that when $x\in\mathbb{R}$, we have
$P_{t}(x)=x\sum_{a,b\geqslant 0}\frac{(-tx)^{a}}{a!}\cdot\frac{(-tx)^{b}}{b!}=xe^{-tx}\cdot e^{-tx}=xe^{-2tx}.$
Hence the eigenvalues of $\mathbf{M}_{t}$ is exactly $P_{t}(\widehat{\lambda}_{j})$. Since
@@
Then by Lemma C.18, we can derive an upper bound. Since $s\leqslant 1$,
@@
$\eqsim\frac{1}{t}\sum_{j=1}^{k}j^{-1+\beta(1-s)}+\sum_{j=k+1}^{N}j^{-1-s\beta}\eqsim\frac{k^{\beta(1-s)}}{t}+k^{-s\beta}.$
@@
$\mathbf{u}_{0}^{\top}\mathbf{A}^{\top}_{t}\mathbf{H}\mathbf{A}_{t}\mathbf{u}_{0}\lesssim\max\{t^{-s},M^{-s\beta}\}\eqsim t^{-s}+M^{-s\beta}.$
Clearly $\mathbf{u}_{0}^{\top}\mathbf{A}^{\top}_{t}\mathbf{H}\mathbf{A}_{t}\mathbf{u}_{0}\lesssim 1$, therefore by Lemma C.5,
$\mathbf{u}_{0}^{\top}\mathbf{A}^{\top}_{t}\mathbf{H}\mathbf{A}_{t}\mathbf{u}_{0}\lesssim\min\{t^{-s},1\}+M^{-s\beta}\eqsim\widehat{e}_{\infty}(t)+M^{-s\beta}\eqsim\widehat{e}_{M}(t)+M^{-s\beta}.$
@@
For $s>0$, it holds with probability at least $1-e^{-\Omega(M)}$ that
$\mathbf{u}_{0}^{\top}\mathbf{A}^{\top}_{t}\mathbf{H}\mathbf{A}_{t}\mathbf{u}_{0}\gtrsim\max\{\widehat{e}_{M}(t),M^{-s\beta}\},$
where $\widehat{e}_{M}(t):=\sum_{j=1}^{M}\widehat{\lambda}_{j}|\theta_{j}^{*}|^{2}e^{-2t\widehat{\lambda}_{j}}$.
@@
where the last inequality is by Von Neumann’s trace inequality: For two symmetric matrices $\mathbf{A}$, $\mathbf{B}\in\mathbb{R}^{d\times d}$ with eigenvalues $a_{1}\geqslant a_{2}\geqslant\dots\geqslant a_{d}$ and $b_{1}\geqslant b_{2}\geqslant\dots\geqslant b_{d}$, we have
$\mathrm{tr}(\mathbf{A}\mathbf{B})\geqslant\sum_{i=1}^{d}a_{i}b_{d+1-i}.$
Note that $\mathbf{M}_{t}=P_{t}(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})$, we then get
@@
$=\sum_{i=1}^{M}e^{-2t\widehat{\lambda}_{i}}\widehat{\lambda}_{i}^{-1}\mu_{i}\left(\mathbf{W}\mathbf{H}\mathbf{u}_{0}\mathbf{u}_{0}^{\top}\mathbf{H}\mathbf{W}^{\top}\right).$
Note that $\mathbf{u}_{0}=\bm{\theta}^{*}$, by Assumption 2.3 and 2.5,
$\mathbf{u}_{0}^{\top}\mathbf{A}^{\top}_{t}\mathbf{H}\mathbf{A}_{t}\mathbf{u}_{0}$ $\geqslant\sum_{i=1}^{M}e^{-2t\widehat{\lambda}_{i}}\widehat{\lambda}_{i}^{-1}\mu_{i}\left(\mathbf{W}\mathbf{H}\mathbf{u}_{0}\mathbf{u}_{0}^{\top}\mathbf{H}\mathbf{W}^{\top}\right)$
$\eqsim\sum_{i=1}^{M}e^{-2t\widehat{\lambda}_{i}}\widehat{\lambda}_{i}^{-1}i^{-1-\beta-s\beta}$
$\eqsim\sum_{i=1}^{M}e^{-2t\widehat{\lambda}_{i}}\widehat{\lambda}_{i}|\theta^{*}_{j}|^{2}=\widehat{e}_{M}(t).$
Here in the second line we used Lemma C.12 for matrix $\mathbf{H}\mathbf{u}_{0}\mathbf{u}_{0}^{\top}\mathbf{H}$ as
$\mu_{j}(\mathbf{H}\mathbf{u}_{0}\mathbf{u}_{0}^{\top}\mathbf{H})=\lambda_{j}^{2}|\theta_{j}^{*}|^{2}\eqsim j^{-1-\beta-s\beta}\implies\mu_{j}(\mathbf{W}\mathbf{H}\mathbf{u}_{0}\mathbf{u}_{0}^{\top}\mathbf{H}\mathbf{W}^{\top})\eqsim j^{-1-\beta-s\beta},$
and the third line follows from again from Lemma C.12 that $\widehat{\lambda}_{j}\eqsim j^{-\beta}$.
@@
$\mathbf{u}_{0}^{\top}\mathbf{A}^{\top}_{t}\mathbf{H}\mathbf{A}_{t}\mathbf{u}_{0}\geqslant\|(\textbf{I}-\mathbf{H}^{\frac{1}{2}}\mathbf{W}^{\top}(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})^{-1}\mathbf{W}\mathbf{H}^{\frac{1}{2}})\mathbf{H}^{\frac{1}{2}}\mathbf{u}_{0}\|^{2}=:T_{3},$